# Continuous emission ultrasound: a new paradigm to ultrafast ultrasound imaging

### v1

@@ Abstract
Current imaging techniques in echography rely on the pulse-echo (PE) paradigm which provides a straight-forward access to the in-depth structure of tissues. They inherently face two major challenges: the limitation of the pulse repetition frequency, directly linked to the imaging framerate, and, due to the emission scheme, their blindness to the phenomena that happen in the medium during the majority of the acquisition time. To overcome these limitations, we propose a new paradigm for ultrasound imaging, denoted by continuous emission ultrasound imaging (CUEI) [1], for a single input single output (SISO) device.
@@ 2.2 Received signal model
The proposed CEUI framework aims at going beyond this consideration, by taking account of a continuous emission interacting with a moving medium. Here is assumed a bi-element ultrasound probe with an emitter at position $\mathbf{p}_{E}$ and a receiver at position $\mathbf{p}_{R}$ with no specific geometry. We consider an imaged medium formed by $N_{S}$ individual point scatterers, indexed by $k$, of respective echogenicity $A_{k}(t)$ and position $\mathbf{p}_{S}^{k}(t)$. The proposed model is mainly based on the time-of-flight of the emitted waveform $x(t)$ between a given scatterer of the medium and a reception element. Note that, given that the scatterers can move during the wave propagation, their time-of-flights can evolve in time.
@@
The proposed model and its implementation are illustrated in Figure 2, for an arbitrary chosen reception time denoted by $t_{R}$. The time-of-flight $T_{TOF}^{k}(t_{E})$ represents the time the ultrasound wave emitted at time $t_{E}$ needs to travel from the emitter to the reflector $k$ at its specific location at the time of interraction, and back to receiving element. It is defined by:
$T_{TOF}^{k}(t_{R})=T_{i\rightarrow k}(t_{R})+T_{k\rightarrow j}(t_{R}),$ (5)
@@
The set defined in (8) characterizes the set scatterers that contributes to the backscattered signal at time $t_{R}$ and describes, for the given reception instant $t_{R}$, which portion of the emission is reflected and the backscattering instant $t_{S}^{k}$ which sets the roundtrip duration of $x(t_{E}^{k})$. The model described in (7) is used to model the step B in Figure 1.
@@ 2.4 Decoding
The previous subsections showed (i) the general continuous emission-reception scheme, that allows the extraction, from continuous emitted and received signals, of $x_{w}(t)$ and $y_{w}(t)$ corresponding to the medium state at a given time $t$, (ii) the excitation waveform used, and (iii) the proposed model of the continuously received signal related to the imaged medium. This subsection explains the decoding step, i.e., how $y_{w}(t)$ is combined with $x_{w}(t)$ to mitigate the effect of the emitted signal. More precisely, matched and mismatched filters are investigated [28], as described hereafter.
@@
The discretized signal $\mathbf{s}_{w}$ is equal to the noise free backscattered echoes exclusively generated by the interaction between the tissues and the signal of reference $\mathbf{x}_{w}$. On the other hand, $\mathbf{v}_{w}$ contains all sources of stochastic noise, plus the echoes generated by the emitted waveforms portions of $\mathbf{x}$ apart from $\mathbf{x}_{w}$. These backscattering are considered as interference because they harm the estimation of the medium which interacts exclusively with $\mathbf{x}_{w}$. Following the model described in subsection 2.3, $\mathbf{s}_{w}$ can be expressed, for a SISO system, as a function of the scatterers forming the medium:
@@
where $\mathbf{I}_{w}$ is the discretisation of a single 1D frame previously noted $I_{w}(r)$ and $\star$ is the cross-correlation operator.
@@ 2.4.2 ISLR mismatched filter
$\mbox{ISLR}\;(\mathbf{c}_{w})=\frac{\mathbf{c}_{w}^{H}\;\mathbf{F}\;\mathbf{c}_{w}}{\mathbf{c}_{w}^{H}\;\overline{\mathbf{F}}\;\mathbf{c}_{w}}\;,\mbox{ with}\mathbf{c}_{w}=\mathbf{\Lambda}_{K}\mathbf{h}_{w},$ (15)
@@ 3 Results
This section will introduce two simulated cases to assess the advantages of CEUI in comparison to pulse-echo, as discussed hereafter. Additionally, Subsection 3.1 will provide an in-depth examination of the simulation configuration and the tuning of hyperparameters within the CEUI framework.
@@ 3.1 Description of the data and simulation parameters
Moreover, an additive white noise is added to the recorded echoes in $\mathbf{y}$ with a SNR set at $10$ dB over the bandwidth of the ultrasound probe.
To generate the recorded RF signal $y(t)$, the model outlined in the subsection 2.3 is employed. The use of this model enables to evaluate the capability of the CEUI system to monitor a dynamic medium with regards to the position and the echogenicity of the scatterers. The spatial dynamic of the medium will cause a non-stationary Doppler effect on the emitted waveform taken into account by the model in (7) and quantified on the following Figure 3. The measured distortions, in orange, prove the spatial dynamic of scatterers is accurately modelled.
@@
The parameters of the CEUI framework are $N_{E}=251$ samples ($\sim 8\mu s$), a step of 21 samples ($\sim 0.7\mu s$) between successive windows, and therefore a slow time frequency $f_{img}$ as high as $1.8$MHz can be achieved. The imaged interval is set between 00 and $4cm$ which leads to $N_{R}\sim 1560$ samples (around $60\mu s$). The raw emitting signal is set as presented in the subsection 2.3.
Figure 3: Doppler effect engendered by a moving scatterer. A monochromatic sinus wave at $f_{c}=5$MHz is sent to a medium containing a single scatterer moving axially at different constant speeds $v_{k}$. The theoretical Doppler frequency shift in blue, is given by $f_{D}=f_{c}\cdot\frac{v_{k}}{c}$, with $v_{k}$ the scatterer axial velocity. The estimated $f_{D}$ in orange, is obtained conducting multiple simulation, with our model, of a scatterer for each different speed $v_{k}$ and to quantify the frequency shift on recorded echoes.
As a reference, M-mode images obtained using a CEUI approach will be compared to M-mode images reconstructed using a PE approach. To do so, a set of pulse emission of identical 13-bits Barker codded emission [30] insonifies the medium. The latter are temporally spaced of $\frac{N_{R}}{f_{s}}\sim 60\mu s$, the pulse repetition interval, providing a framerate of $f_{img,PE}=18kHz$ considering too, an imaging range from 0 to $4$cm and an homogeneous wave velocity throughout the medium $c=1540\;m.s^{-1}$. This is a 100 times smaller compared to CEUI slow time frequency. The associated decoding method for PE is matched filter.
@@ 3.2 Fast dynamic event imaging
Figure 4: M-Mode imaging of an axially oscillating scatterer at $12$kHz around $30$mm. PE at a framerate of $19$kHz is performed on the left as described earlier, and CEUI at a slow time frequency of $1.4$MHz is used in the middle and on the right with respectively MF and wide misMF as decoding filters.
@@ 3.3 Quick appearing event imaging
The simulation illustrated on Figure 5 highlights the main contribution of the continuous insonification while using the CEUI approach: because the continuous emitted waveform and the echogenic structure in the field of view are interacting without interruption, our imaging system is capable to capture short phenomenons lasting only a dozen of $\mu$s. The same probe and post-processing settings as described previously are performed for a simulated medium containing a spatially static blinking scatterer at $30$mm depth.
@@
The upper graph on Figure 5 displays the ground truth blinks of the scatterer during periods in blue. The horizontal grey line represents the scatterer depth at 30mm, while the green areas are the wavefronts of the successive pulse emissions of PE approach. When a green beam meets the blue area at the scatterer depth, it means the current pulse is at least partially backscattered which permits the detection: this is the case for the $5^{th}$ and $9^{th}$ scatterer apparitions. No backscattering is produced during other apparitions of the scatterer. The difference of amplitude between the two detected blinks depends on the proportion of the emitted pulse which interacts with the scatterer.
The improper blinking identification using PE leads to a bad interpretation of the events occurring within the medium. Around $200\mu$s, multiple instances of blinking (5 and 6) are confused with a single occurrence, or conversely, situations where an actual appearance takes place go unnoticed as shown for blinks 1, 2, 3, 4, 6, 7, 8 and 10.
While PE reveals limitations to detect brief events with regards to the pulse repetition interval, the two lower M-mode images on Figure 5, obtained using CEUI, depict well all the 10 blinks. RF data can record echoes regardless of the duration of the scatterer’s appearance. The accurate identification of these phenomena then hinges on their duration and the intervals between them at the same depth, both intricately linked to the tuning of the parameter $N_{E}$.
@@
Figure 5: M-Mode images of a spatially static blinking scatterer located at $30$mm depth. The blinking instants are randomly distributed during the whole acquisition. The first graph displays, in blue the periods when the scatterer is echogeneous, in green the ranges insonified by the emitted waveform by PE, and in gray, the horizontal line is located at the depth of the scatterer. The first M-mode (top to bottom) image results from a PE approach at $19$kHz, while the second and third ones use a CEUI approach at $1.4$MHz with respectively the MF and the misMF as decoding techniques.

### v2

@@ Abstract
Current imaging techniques in echography rely on the pulse-echo (PE) paradigm which provides a straight-forward access to the in-depth structure of tissues. They inherently face two major challenges: the limitation of the pulse repetition frequency, directly linked to the imaging framerate, and, due to the emission scheme, their blindness to the phenomena that happen in the medium during the majority of the acquisition time. To overcome these limitations, we propose a new paradigm for ultrasound imaging, denoted by continuous emission ultrasound imaging (CEUI) [1], for a single input single output (SISO) device.
@@ 2.2 Received signal model
The proposed CEUI framework aims at going beyond this consideration, by taking account of a continuous emission interacting with a moving medium. Here is assumed a bi-element ultrasound probe with an emitter at position $\mathbf{p}_{E}$ and a receiver at position $\mathbf{p}_{R}$ with no specific geometry. We consider an imaged medium formed by $N_{S}$ individual point scatterers, indexed by $k$, of respective echogenicity $A_{k}(t)$ and position $\mathbf{p}_{S}^{k}(t)$. The latter parameter, in addition to the emitter and receiver element positions, enables to compute the most important feature of our model: the non stationary time-of-flight (TOF) $T_{TOF}^{k}(t)$ of each scatterer through the acquisition time. Indeed, it is crucial to note that scatterers motion is modeled during the propagation of the emitted wave through the medium.
@@
The proposed model and its implementation are illustrated in Figure 2, for an arbitrary chosen reception time denoted by $t_{R}$. The time-of-flight $T_{TOF}^{k}(t_{E})$ represents the time the ultrasound wave emitted at time $t_{E}$ needs to travel from the emitter to the reflector $k$ at its specific location at the time of interaction, and back to receiving element. It is defined by:
$T_{TOF}^{k}(t_{R})=T_{E\rightarrow k}(t_{R})+T_{k\rightarrow R}(t_{R}),$ (5)
@@
The set defined in (8) characterizes the set scatterers that contributes to the backscattered signal at time $t_{R}$ and describes, for the given reception instant $t_{R}$, which portion of the emission is reflected and the backscattering instant $t_{S}^{k}$ which sets the roundtrip duration of $x(t_{E}^{k})$. The model described in (7) is used to model the step B in Figure 1. Its implementation, assuming a linear interpolation of non integer backscattering and emitting times, is resumed in the following pseudo-code.
Algorithm 1 RF signal generation for a dynamic medium assuming no attenuation
0:  $(\mathbf{p}_{S}^{k})_{k}$, $(\mathbf{A}_{k})_{k}$, $\mathbf{p}_{E}$, $\mathbf{p}_{R}$, $c$, $\mathbf{e}$,
0:  $\mathbf{y}$
  $\mathbf{y}=\mathbf{0}$
  for $k\in\llbracket 1,\;N_{S}\rrbracket$ do
     $\mathbf{y_{k}}=\mathbf{0}$
     for $n_{R}\in\llbracket 1:N_{echo}\rrbracket$ do
        $n_{S}^{k}$ s.t. $n_{R}=\frac{R_{k\rightarrow R}(n_{S}^{k})}{c\cdot T_{s}}+n_{S}^{k}$
        $n_{E}^{k}\leftarrow n_{S}^{k}-\frac{R_{E\rightarrow k}(n_{S}^{k})}{c\cdot T_{s}}$
        $\alpha\leftarrow n_{E}^{k}-\lfloor n_{E}^{k}\rfloor$
        $\beta\leftarrow n_{S}^{k}-\lfloor n_{S}^{k}\rfloor$
        $\mathbf{y}{[n_{R}]}\leftarrow\Bigl{(}(1-\beta)\cdot\mathbf{A}_{k}{\bigl{[}\lfloor n_{S}^{k}\rfloor\bigr{]}}+\beta\cdot\mathbf{A}_{k}{\bigl{[}\lfloor n_{S}^{k}\rfloor+1\bigr{]}}\Bigr{)}\cdot\Bigl{(}(1-\alpha)\cdot\mathbf{x}{\bigl{[}\lfloor n_{E}^{k}\rfloor\bigr{]}}+\alpha\cdot\mathbf{x}{\bigl{[}\lfloor n_{E}^{k}\rfloor+1\bigr{]}}\Bigr{)}$
     end for
     $\mathbf{y}\leftarrow\mathbf{y}+\mathbf{y}_{k}$
  end for
@@ 2.4 Decoding
The previous subsections showed (i) the general continuous emission-reception scheme, that allows the extraction, from continuous emitted and received signals, of $x_{w}(t)$ and $y_{w}(t)$ corresponding to the medium state at a given time $t$, (ii) the excitation waveform used, and (iii) the proposed model of the continuously received signal related to the imaged medium. This subsection explains the decoding step, i.e., how $y_{w}(t)$ is combined with $x_{w}(t)$ to mitigate the effect of the emitted signal. More precisely, matched and mismatched filters are investigated [28], as described hereafter. In contrast to PE, where the energy of the recorded signal is mostly concentrated at times directly associated to scatterer positions, CEUI produces echoes throughout the whole acquisition time. Therefore, a decoding and unmixing process is needed to reconstruct spatial structures.
@@
The discretized signal $\mathbf{s}_{w}$ is equal to the noise free backscattered echoes exclusively generated by the interaction between the tissues and the signal of reference $\mathbf{x}_{w}$. On the other hand, $\mathbf{v}_{w}$ contains all sources of stochastic noise, plus the echoes generated by the emitted waveforms portions of $\mathbf{x}$ apart from $\mathbf{x}_{w}$. These backscattering are considered as interference because they harm the estimation of the medium which interacts exclusively with $\mathbf{x}_{w}$. These interference, that may harm the resulting image contrast, are mitigated by the significant higher energy contained by $\mathbf{x}_{w}$ compared to the one carried by a single emitted pulse in the PE paradigm. Assuming both aforementioned signals have approximately the same power, the energy ratio equals to the ratio between their temporal length. Following the model described in subsection 2.3, $\mathbf{s}_{w}$ can be expressed, for a SISO system, as a function of the scatterers forming the medium:
@@
where $\mathbf{I}_{w}$ is the discretization of a single 1D frame previously noted $I_{w}(z)$ and $\star$ is the cross-correlation operator.
@@ 2.4.2 ISLR mismatched filter
$\mbox{ISLR}\;(\mathbf{c}_{w})=\frac{\mathbf{c}_{w}^{H}\;\mathbf{F}\;\mathbf{c}_{w}}{\mathbf{c}_{w}^{H}\;\overline{\mathbf{F}}\;\mathbf{c}_{w}}\;,\mbox{ with}\mathbf{c}_{w}=\mathbf{\Lambda}_{K}\mathbf{x}_{w},$ (15)
@@ 3 Results
This section studies four simulation cases. The first two assess the advantages of CEUI in comparison to PE to image quick (see Subsection 3.3) and fast (see Subsection 3.2) events, as discussed hereafter. Then, CEUI is evaluated within more realistic simulations, with regards to the sparsity of the medium (see Subsection 3.4). Finally, a study on the potential benefits of CEUI to characterize slower phenomenon is investigated, especially with regards to the investigated depth of the imaging system (see Subsection 3.5). Additionally, Subsection 3.1 provides an in-depth examination of the simulation configuration and hyperparameter tuning within the CEUI framework.
@@ 3.1 Description of the data and simulation parameters
Moreover, an additive white noise is added to the recorded echoes in $\mathbf{y}$ with a SNR set at $10$ dB. The latter is computed as the power ratio between the emitted signal $\mathbf{x}$ and the noise over the bandwidth of the ultrasound probe.
To generate the recorded RF signal $y(t)$, the model outlined in the subsection 2.3 is employed. The use of this model enables to evaluate the capability of the CEUI system to monitor a dynamic medium with regards to the position and the echogenicity of the scatterers. The spatial dynamic of the medium will cause a non-stationary Doppler effect on the emitted waveform taken into account by the model in (7) and quantified on the following Figure 3.
Figure 3: Doppler effect engendered by a moving scatterer. A monochromatic sinus wave at $f_{c}=5$MHz is sent to a medium containing a single scatterer moving axially at different constant speeds $v_{k}$. The theoretical Doppler frequency shift in blue, is given by $f_{D}=f_{c}\cdot\frac{v_{k}}{c}$, with $v_{k}$ the scatterer radial velocity. The estimated $f_{D}$ in orange, is obtained conducting multiple simulation, with our model, of a scatterer for each different speed $v_{k}$ and to quantify the frequency shift on recorded echoes.
To estimate each distortion, a simulated scatterer at a given radial velocity $v_{k}$ is modelled and insonified by a monochromatic emission at $f_{c}$. The frequency Doppler shift on the recorded echoes is then estimated for each radial velocity. Comparing the measured distortions, in orange, to theoretical values, in blue, the spatial dynamic of scatterers is accurately modelled in our simulations.
@@
The parameters of the CEUI framework are defined as follows: the reference signal, $N_{E}$, is set to 251 samples ($\sim 8\mu s$) for Subsections 3.2 and 3.3, 2001 samples ($\sim 65\mu s$) for Subsection 3.4 and 4501 samples ($\sim 145\mu s$) in Subsection 3.5. The latter parameter increases when imaged phenomenon are low and vice versa. A step of 21 samples ($\sim 0.7\mu s$) between successive windows is used. Therefore a slow time frequency $f_{img}$ as high as $1.8$MHz can be achieved. The imaged interval is set between 00 and $4$cm which leads to $N_{R}\sim N_{E}+1560$ samples.
The M-Mode images in Figures 4 and 5 are shown on a narrow depth range to identify more accurately the difference of performance between PE and CEUI. Nevertheless, note that CEUI slow time frequency is not constrained by the imaging range, compared to PE one. The raw emitting signal is set as presented in the subsection 2.3.
As a reference, M-mode images obtained using a CEUI approach will be compared to M-mode images reconstructed using a PE approach. To do so, a set of pulse emission of identical 13-bits Barker codded emission [30] insonifies the medium. The latter are temporally spaced of $\frac{2\cdot R_{max}}{c}$, the pulse repetition interval, providing a framerate of $f_{img,PE}=19kHz$ considering too, an imaging range from 0 to $4$cm and an homogeneous wave velocity throughout the medium $c=1540\;m.s^{-1}$. This is a 100 times smaller compared to CEUI slow time frequency. The associated decoding method for PE is matched filter.
@@ 3.2 Fast dynamic event imaging
Figure 4: M-Mode imaging of an axially oscillating scatterer at $12$kHz around $30$mm. The output image using PE at a slow time frequency of $19$kHz is displayed, while CEUI at a slow time frequency of $1.4$MHz is showed in the middle, respectively below with MF, respectively mMF as decoding methods.
@@ 3.3 Quick appearing event imaging
The simulation illustrated on Figure 5 highlights the main contribution of the continuous insonification while using the CEUI approach: because the continuous emitted waveform and the echogenic structure in the field of view are interacting without interruption, our imaging system is capable to capture short phenomenons lasting only a dozen of $\mu$s. The same probe and post-processing settings as described previously are performed for a simulated medium containing a spatially static blinking scatterer at $30$mm depth with a randomly distributed echogenicity at each apparition.
@@
The upper graph on Figure 5 displays the ground truth blinks of the scatterer during periods in blue. The horizontal grey line represents the scatterer depth at 30mm, while the green areas are the wavefronts of the successive pulse emissions of PE approach. When a green beam meets the blue area at the scatterer depth, it means the current pulse is at least partially backscattered which permits the detection: this is the case for the $5^{th}$ and $9^{th}$ scatterer apparitions. No backscattering is produced during other apparitions of the scatterer. The difference of amplitude between the two detected blinks depends on the proportion of the emitted pulse which interacts with the scatterer. The improper blinking identification using PE leads to a bad interpretation of the events occurring within the medium. Around $200\mu$s, multiple instances of blinking (7, 16 and 19) are confused with a single occurrence, or conversely, situations where an actual appearance takes place go unnoticed (every other blinks). It is concretely shown by the red signal of Figure 5 fourth plot where only 3 echoes are recorded out of 20 potential ones. Under PE paradigm, blinks identification relies on multiple factors which are: the apparition frequency, their duration and their temporal localization. However, even when apparitions are detected by PE, the proportion of pulses interacting with the scatterers can introduce a bias in the relative echogenecities imaged in the M-Mode.
@@
Figure 5: M-Mode images of a spatially static blinking scatterer located at $30$mm depth. The blinking instants are randomly distributed during the whole acquisition. The echogenecity $A_{k}$ is also uniformly random distributed between 0.5 and 1. The first graph displays, in blue the periods when the scatterer is echogeneous, in green the ranges insonified by the emitted waveform by PE, and in gray, the horizontal line is located at the depth of the scatterer. The first M-mode (top to bottom) image results from a PE approach at $19$kHz, while the second and third ones use a CEUI approach at $1.4$MHz with respectively the MF and the misMF as decoding techniques. The fourth figure plots the ground truth M-Mode line at 30mm depth and its estimation in pink, respectively red, using CEUI respectively PE.
For instance, the 7th blink interacts only with the tail end of the pulse, resulting in a falsely lower representation compared to the 15th blink.
While PE reveals limitations to detect brief events with regards to the pulse repetition interval, the two lower M-mode images on Figure 5, obtained using CEUI, depict well all the 20 blinks. RF data can record echoes regardless of the duration of the scatterer’s appearance. The accurate identification of these phenomena then hinges on their duration and the intervals between them at the same depth, both intricately linked to the tuning of the parameter $N_{E}$. In addition, the intensity contrast difference between blinks is generally well depicted on both CEUI M-Mode images thanks to a complete interaction during each apparition period.
3.4 Cyst imaging with speckle
Due to the continuous insonification of the medium, in a more realistic scenario, the emitted waveform propagates through a medium, generating low-energy echoes that contribute to the speckle in the image. The following experiment aims at evaluating the capacity of CEUI to identify a cyst within a material background whereas the energy of interference signal $\mathbf{v}_{w}$ is significantly more important than in experiments led in subsections 3.2 and 3.3. This is due to the increased number of echogenic cyst scatterers plus the presence of a background. The simulated M-Mode images are represented in Figure 6.
Note that $N_{E}$ was increased in this experiment from 251 samples to 2001 samples, to increase decoding performances in the presence of a slower moving big structure. However, it does not impact the slow time frequency of the resulting M-Mode which remains at $1.4$ MHz. Thanks to this frequency much higher than the bound imposed by Shannon theorem, it is even possible to perform temporal compounding using neighbor 1D-lines to improve image quality. For illustration purpose, a hamming window was applied in the summation of the 1D-lines to preserve temporal resolution.
Figure 6: M-Mode images of dilated and compressed cyst surrounded by a background. The cyst is modelled by 50 scatterers axially spaced, in average, by $\lambda$ with an average echogenecity of 1. The background contains 10 scatterers randomly set every $\lambda$ in depth with an echogenecity defined by a Rayleigh distribution centered in 0.05. The total number of background scatterers reaches 950.
Qualitatively, both decoding methods applied using CEUI enable clear identification of the central dilating cyst. The boundaries of the object are distinctly depicted. The deeper the scatterers in the cyst, the less they move which therefore, increases decoding performances of their echoes. For both PE and CEUI, the maximal dilatation is better depicted because the medium reaches a zero velocity state. Quantitatively, the contrast between the cyst and the background increases the SNR of $3$ dB, (respectively remains the same) using mMF (respectively MF) as CEUI decoder compared to PE.
3.5 Contrast improvement in a medium with attenuation
The current subsection investigates the pertinence of the use of CEUI to image a slower medium. To do so the peak signal-to-noise ratio (PSNR), as indicator of the image contrast, will be locally quantified on M-Mode images of a linearly attenuated medium defined by 6 static scatterers. The attenuation is modelled, similarly to ultrasound simulators such as Field II [31], by:
$A_{att}^{k}(R_{E\rightarrow k},R_{k\rightarrow R})=exp\big{(}-\alpha\cdot f_{c}\cdot(R_{E\rightarrow k}+R_{k\rightarrow R})\big{)}\$ (19)
with $\alpha=1.5$ dB/MHz/cm. The scatterers are placed at depths of 30, 45, 60, 75, 90 and 105 mm, with the same constant echogenecity through the acquisition time.
Figure 7: Regions of interest on M-Mode images built of a medium composed of 6 static scatterers with an equal static echogenecity during all the acquisition time. A linear attenuation due to the propagation of the emission through the medium is incorporated with a coefficient $\alpha=1.5$dB/cm/MHz. From left to right, the methods used are CEUI with MF decoding, CEUI with mMF decoding and finally PE with MF based pulse compression. The last column images show the average profile obtained using each method at each depth.
[…]
$N_{E}\leq\frac{\pi\cdot\Delta z\cdot f_{S}}{v_{M}}\sim 4501\mbox{ samples}$ (20)
The resulting M-Mode images are regrouped in Figure 7. One can note that PE generates higher main-lobe peaks, but with significantly higher sidelobes. Computing the average over slow time of the PSNR on each 1D-line around each scatterer depth, quantifies the image contrast. CEUI-mMF improves the PSNR by 3 to 5 dB depending on the image depth. Consequently, still based on the assumption of a linearly attenuated medium, and considering slower phenomenon to enable a wider reference signal, improves the investigated depth by 2 cm without compromising the temporal resolution of the acquisition.